OpenStudy (anonymous):

|x - 10| - 16 = -9 absolute value help!!

OpenStudy (anonymous):

The first step in these problems is to get the absolute value y itself on one side. So do this first

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

|x-10|=7 means 2 things (x-10)=7 and -(x-10)=7 Because the absolute value creates 2 answers

OpenStudy (anonymous):

You need to solve both equations to get both values of x

OpenStudy (anonymous):

@ChmE okay thanks!!